# Interference Among First-Price Pacing Equilibria: A Bias and Variance Analysis

### v1

@@ 4.1. A Debiased Surrogate for Pacing Multipliers
If we view $\beta^{*}_{\alpha}$ as a function of the level of contamination $\alpha$, then one can imagine that under sufficient regularity conditions, the pacing multipliers in the perfectly separated market, $\beta^{*}_{0}$, can be approximated by some form of Taylor expansion of $\alpha\mapsto\beta^{*}_{\alpha}$ at $\alpha$. This can be made rigorous by the notion of directional derivatives. We define
(5) $\mathop{}\!\mathrm{d}\beta^{*}(\alpha)=\lim_{\delta\to 0^{+}}\frac{\beta^{*}_{\alpha-\delta}-\beta^{*}_{\alpha}}{\delta}\;.$
@@
if the limit exists. We will show that $\beta^{*}_{\alpha}+\alpha\mathop{}\!\mathrm{d}\beta^{*}(\alpha)$ serves as a good approximation to $\beta^{*}$.
@@
We now show that the directional derivative $\mathop{}\!\mathrm{d}\beta^{*}(\alpha)$ has a closed-form expression under certain regularity conditions (the conditions are given in Thm. 1; the full proof is given in the appendix). We need a few notations for this expression. Define $\Delta_{\alpha}=\int\nabla f(\theta,\beta^{*}_{\alpha})(s-s^{\prime})\mathop{}\!\mathrm{d}\theta.$ Let ${\mathscr{H}}_{\alpha}=\nabla_{\beta\beta}^{2}\int F(\theta,\beta^{*}_{\alpha})s_{\alpha}\mathop{}\!\mathrm{d}\theta$ be the Hessian matrix in the market ${\mathscr{M}}_{\alpha}$ and $P_{\alpha}={\operatorname*{Diag}}(1(\beta^{*}_{\alpha,i}<1))$. Then, under regularity conditions,
(6) $\mathop{}\!\mathrm{d}\beta^{*}(\alpha)=-(P_{\alpha}{\mathscr{H}}_{\alpha}P_{\alpha})^{\dagger}\Delta_{\alpha}\;.$
We present a heuristic argument for Eq. 6 here; for a rigorous treatment we need a perturbation theory result which is given in Lemma 1 in the appendix. Define the contaminated EG objective $H_{\alpha}(\beta)=\int F(\beta,\theta)s_{\alpha}\mathop{}\!\mathrm{d}\theta$. Then ${\mathscr{H}}_{\alpha}=\nabla^{2}H_{\alpha}(\beta^{*}_{\alpha})$. The map $(\beta,\alpha)\mapsto H_{\alpha}(\beta)$ has gradient $[\nabla H(\beta^{*}),\int F(\beta^{*}_{\alpha},\theta)(s^{\prime}-s)\mathop{}\!\mathrm{d}\theta]^{\scriptscriptstyle\mathsf{T}}$ and Hessian $[{\mathscr{H}}_{\alpha},-\Delta_{\alpha};-\Delta_{\alpha}^{\scriptscriptstyle\mathsf{T}},0]$ at $(\beta^{*}_{\alpha},\alpha)$. Then we have the following quadratic approximation of $H_{\alpha-\delta}(\beta)$
$H_{\alpha-\delta}(\beta)\approx H_{\alpha}(\beta^{*}_{\alpha})+\begin{pmatrix}\beta-\beta^{*}_{\alpha}\\ -\delta\end{pmatrix}^{\scriptscriptstyle\mathsf{T}}\begin{pmatrix}\nabla H_{\alpha}(\beta^{*}_{\alpha})\\ \int F(\beta^{*}_{\alpha},\theta)(s^{\prime}-s)\mathop{}\!\mathrm{d}\theta\end{pmatrix}+\frac{1}{2}\begin{pmatrix}\beta-\beta^{*}\\ -\delta\end{pmatrix}^{\scriptscriptstyle\mathsf{T}}\begin{pmatrix}{\mathscr{H}}_{\alpha}&-\Delta_{\alpha}\\ -\Delta_{\alpha}^{\scriptscriptstyle\mathsf{T}}&0\end{pmatrix}\begin{pmatrix}\beta-\beta^{*}\\ -\delta\end{pmatrix}\;.$
And so
$\beta^{*}_{\alpha-\delta}=\operatorname*{arg\,min}_{\beta\in(0,1]^{n}}H_{\alpha-\delta}(\beta)\approx\operatorname*{arg\,min}_{\beta\in(0,1]^{n}}(\nabla H_{\alpha}(\beta^{*}_{\alpha})+\delta\Delta_{\alpha})^{\scriptscriptstyle\mathsf{T}}\beta+\tfrac{1}{2}(\beta-\beta^{*}_{\alpha})^{\scriptscriptstyle\mathsf{T}}{\mathscr{H}}_{\alpha}(\beta-\beta^{*}_{\alpha})\;,$
where we dropped terms that do not involve $\beta$.
Next we need two observations. For $\delta$ small enough, we have $\beta^{*}_{\alpha-\delta,i}=1$ if $\beta^{*}_{\alpha,i}=1$. Intuitively, buyers in the market ${\mathscr{M}}_{\alpha}$ with $\beta^{*}_{\alpha,i}=1$ have some nonzero leftover budget, which acts as a buffer to small changes in supply, and thus their pacing multipliers remain one. Next, we also note $\beta\mapsto\nabla H_{\alpha}(\beta^{*}_{\alpha})^{\scriptscriptstyle\mathsf{T}}\beta$ is constant on the set $\{\beta:{\beta_{i}}=1\text{ if }\beta^{*}_{\alpha,i}=1\}$ by complementary slackness. Then we further simplify and obtain
$\beta^{*}_{\alpha-\delta}\approx\operatorname*{arg\,min}_{{\beta_{i}}=1\text{ if }\beta^{*}_{\alpha,i}=1}\delta\Delta_{\alpha}^{\scriptscriptstyle\mathsf{T}}\beta+\tfrac{1}{2}(\beta-\beta^{*}_{\alpha})^{\scriptscriptstyle\mathsf{T}}{\mathscr{H}}_{\alpha}(\beta-\beta^{*}_{\alpha})\;.$
The right-hand side is just a linearly constrained quadratic optimization problem, which admits a closed-form solution $\beta^{*}_{\alpha}-\delta(P_{\alpha}{\mathscr{H}}_{\alpha}P_{\alpha})^{\dagger}\Delta_{\alpha}$. Now we define a debiased pacing multiplier
(7) $\widetilde{\beta}^{*}=\beta^{*}_{\alpha}+\alpha\cdot(-(P_{\alpha}{\mathscr{H}}_{\alpha}P_{\alpha})^{\dagger}\Delta_{\alpha})\;.$
@@
Suppose that in the market ${\mathscr{M}}_{0}$ conditions Definition 2 (SMO). and Definition 3 (SCS). hold, and assume that $\beta\mapsto\nabla^{2}\int F(\beta,\theta)s^{\prime}\mathop{}\!\mathrm{d}\theta$ is twice continuously differentiable in a neighborhood of $\beta^{*}$. Then the directional derivative $\mathop{}\!\mathrm{d}\beta^{*}({\cdot})$ is well-defined in a neighborhood of zero, and is given by Eq. 6. Moreover, as $\alpha\downarrow 0$,
$\|\widetilde{\beta}^{*}-\beta^{*}\|_{2}=o(\alpha)\;.$
@@
The proof of Thm. 1 is given in Sec. B.1. Thm. 1 indicates that the debiased surrogate $\widetilde{\beta}^{*}$ removes first-order bias caused by interference. The limit pacing multipliers $\beta_{\alpha}^{*}$ of the contaminated market ${\mathscr{M}}_{\alpha}$ will have bias of order $\beta^{*}_{\alpha}-\beta^{*}=\Theta(\alpha)$. In contrast, Thm. 1 shows that the debiased surrogate only incurs a bias of order $o(\alpha)$.
4.2. The Estimator
@@
In this section we introduce a plug-in estimator for the debiased surrogate $\widetilde{\beta}^{*}$ and introduce a consistency theorem. In next section we discuss constructing confidence intervals.
@@
To estimate $\mathop{}\!\mathrm{d}\beta^{*}(\alpha)$ in Eq. 6 we need estimates of its three components: the Hessian ${\mathscr{H}}_{\alpha}=\nabla_{\beta\beta}^{2}\int F(\theta,\beta^{*}_{\alpha})s_{\alpha}\mathop{}\!\mathrm{d}\theta$, the diagonal matrix $P_{\alpha}$ and the vector $\Delta_{\alpha}=\int\nabla f(\theta,\beta^{*}_{\alpha})(s-s^{\prime})\mathop{}\!\mathrm{d}\theta$.
@@
  • The Hessian. For simplicity in our theoretical results, we will simply assume a generic Hessian estimator ^$\widehat{\mathscr{H}}_{\alpha}$ such that for some $\eta_{t}\downarrow 0$ we have ^$\widehat{\mathscr{H}}_{\alpha}-{\mathscr{H}}_{\alpha}=O_{p}(\eta_{t})$. Hong et al. (2015) discuss the estimation of the derivative in detail. Different kinds of statistical guarantees require different rate conditions on $\eta_{t}$; see Theorems 2 and 3. We then introduce two Hessian estimators: one is applicable for general FPPE, while the other requires an extra market regularity condition. The first Hessian estimator is the finite difference method. Let $e_{i},e_{j}$ be basis vectors and $\varepsilon_{t}$ be a step-size. Then the estimator is
    ^$\widehat{\mathscr{H}}_{\alpha}[i,j]=[H_{\alpha,t}(\beta^{\gamma}_{++})-H_{\alpha,t}(\beta^{\gamma}_{+-})-H_{\alpha,t}(\beta^{\gamma}_{-+})+H_{\alpha,t}(\beta^{\gamma}_{--})]/(4\varepsilon_{t}^{2})\;,$
@@
    where $\beta^{\gamma}_{\pm\pm}=\beta^{\gamma}_{\alpha}\pm e_{i}\varepsilon_{t}\pm e_{j}\varepsilon_{t}$, and $H_{\alpha,t}(\beta)=\frac{1}{t}{\sum_{\tau=1}^{t}}F({\theta^{\tau}},\beta)$, with $\{{\theta^{\tau}}\}_{\tau}$ being the items in ^$\widehat{\mathscr{M}}_{\alpha}$. In practice, a diagonal approximation of the Hessian suffices. The second method relies on an additional regularity condition, in which case we derive a simplified formula for the Hessian, thereby enabling a simpler estimation procedure (see Thm. 3).
@@
  • The vector $\Delta_{\alpha}$. Let $g$ be the Radon-Nikodym ratio $g(\theta)=(\mathop{}\!\mathrm{d}(s-s^{\prime})/\mathop{}\!\mathrm{d}s_{\alpha})(\theta)=\frac{1}{1-\alpha}1_{\Theta_{\mathsf{good}}}(\theta)-\frac{1}{\alpha}1_{\Theta_{\mathsf{bad}}}(\theta)$. With the ratio $g$, the true vector $\Delta_{\alpha}$ can be written as $\Delta_{\alpha}=\int g(\theta)\nabla f(\theta,\beta^{*}_{\alpha})s_{\alpha}(\theta)\mathop{}\!\mathrm{d}\theta$, which is easy to estimate given i.i.d. draws from $s_{\alpha}$. In particular, our estimator is then
    $\widehat{\Delta}_{\alpha}=\frac{1}{t}{\sum_{\tau=1}^{t}}g({\theta^{\tau}})\mu^{\tau}\;.$
    Here $\mu^{\tau}=[x_{1}^{\tau}v_{1}^{\tau},\dots,x_{n}^{\tau}v_{n}^{\tau}]^{\scriptscriptstyle\mathsf{T}}$ is a subgradient of $f({\theta^{\tau}},\beta^{\gamma}_{\alpha})$ w.r.t. $\beta$.
@@
  • The diagonal matrix $P_{\alpha}$. Recall $P_{\alpha}={\operatorname*{Diag}}(1(\beta^{*}_{\alpha,i}<1))$. So a natural estimator is $\widehat{P}_{\alpha}={\operatorname*{Diag}}(1(\beta^{\gamma}_{\alpha,i}<1-\iota_{t}))$, where the slackness $\iota_{t}\asymp\frac{1}{\sqrt{t}}$.
@@
With all three components estimated, using $\beta^{\gamma}_{\alpha}$ is the pacing multiplier in the market ^$\widehat{\mathscr{M}}_{\alpha}$ we define the plug-in estimator for $\mathop{}\!\mathrm{d}\beta^{*}(\alpha)$ in Eq. 6 as
(8) $\widehat{\beta}=\beta^{\gamma}_{\alpha}-\alpha\,{\cdot}\,(\widehat{P}_{\alpha}\widehat{\mathscr{H}}_{\alpha}\widehat{P}_{\alpha})^{\dagger}\widehat{\Delta}_{\alpha}\;.$
@@
4.3. Asymptotic Normality and Inference
We present two asymptotic normality results. In the first result, we require a stronger condition on the Hessian error rate $\eta_{t}$. In particular, as will be shown in Thm. 3, the rate condition $\eta_{t}=o(1/\sqrt{t})$ is sufficient for normality. One could use a separate large historical dataset to obtain a good estimate of the Hessian matrix. In the second result, we impose an additional condition on market structure which simplifies the Hessian expression and facilitates efficient Hessian estimation. To describe the additional market structure, we define the gap between the highest and the second-highest bid for an item $\theta$ under pacing $\beta$ by
${\mathsf{{bidgap}}}(\beta,\theta)=\max\{{\beta_{i}}{v_{i}(\theta)}\}-\operatorname{secondmax}\{{\beta_{i}}{v_{i}(\theta)}\}\;,$
where $\operatorname{secondmax}$ is the second-highest entry potentially equal to the highest; e.g., $\operatornamewithlimits{secondmax}([1,1,2])=1$. When there is a tie for an item $\theta$ under pacing $\beta$, we have ${\mathsf{{bidgap}}}(\beta,\theta)=0$. When there is no tie for an item $\theta$, the gap ${\mathsf{{bidgap}}}(\beta,\theta)$ is strictly positive.
For any $g:\Theta\to{\mathbb{R}}$, let ${\mathbb{E}}_{\alpha}[g]=\int gs_{\alpha}\mathop{}\!\mathrm{d}\theta$ and ${\operatorname{Cov}}_{\alpha}(g)={\mathbb{E}}_{\alpha}[(g-{\mathbb{E}}_{\alpha}[g])(g-{\mathbb{E}}_{\alpha}[g])^{\scriptscriptstyle\mathsf{T}}]$. Recall $\widetilde{\beta}^{*}$ is the debiased surrogate in Eq. 7 and $\widehat{\beta}$ is its estimator defined in Eq. 8.
We need to introduce a few more notations to describe the normality results. First, let
$D_{\alpha}=-(P_{\alpha}{\mathscr{H}}_{\alpha}P_{\alpha})^{\dagger}\nabla f({\cdot},\beta^{*}_{\alpha})\;,$
As mentioned previously, the pacing multipliers in the contaminated market converge to the limit counterpart and have the representation
$\sqrt{t}(\beta^{\gamma}_{\alpha}-\beta^{*}_{\alpha})=\frac{1}{\sqrt{t}}{\sum_{\tau=1}^{t}}(D_{\alpha}({\theta^{\tau}})-{\mathbb{E}}_{\alpha}[D_{\alpha}({\theta^{\tau}})])+o_{p}(1)\;.$
In the statistics literature, the function $D_{\alpha}({\cdot})-{\mathbb{E}}[D_{\alpha}]$ is called the influence function (van der Vaart, 2000). For our debiased estimators, we need the following (uncentered) influence functions.
$IF_{1}(\theta)=\big{(}\tfrac{1}{1-\alpha}1_{\Theta_{\mathsf{good}}}(\theta)\big{)}D_{\alpha}(\theta)\;,$
$IF_{2}(\theta)=IF_{1}(\theta)-2\alpha{\operatorname*{Diag}}(\beta^{*}_{\alpha,i}\Delta_{\alpha,i}b_{i}^{{-1}})D_{\alpha}(\theta)\;.$
Theorem 3 (Asymptotic Normality).
Let Definition 3 (SCS). and Definition 2 (SMO). hold in ${\mathscr{M}}_{\alpha}$.
1. It holds that $\widehat{\beta}-\widetilde{\beta}^{*}={Z_{t}}+O_{p}(\eta_{t})+o_{p}(\tfrac{1}{\sqrt{t}}),$ where

### v3

@@ 4.1 A Debiased Surrogate for Pacing Multipliers
If we view ${\bm{\beta}}^{*}_{\alpha}$ as a function of the level of contamination $\alpha$, then one can imagine that under sufficient regularity conditions, the pacing multipliers in the perfectly separated market, ${\bm{\beta}}^{*}_{0}$, can be approximated by some form of Taylor expansion of $\alpha\mapsto{\bm{\beta}}^{*}_{\alpha}$ at $\alpha$. This can be made rigorous by the notion of directional derivatives. We define
$\mathop{}\!\mathrm{d}{\bm{\beta}}^{*}(\alpha)=\lim_{\epsilon\to 0^{+}}\frac{{\bm{\beta}}^{*}_{\alpha-\epsilon}-{\bm{\beta}}^{*}_{\alpha}}{\epsilon}\;.$ (5)
@@
if the limit exists. We will show in Thm. 1 that ${\bm{\beta}}^{*}_{\alpha}+\alpha\mathop{}\!\mathrm{d}{\bm{\beta}}^{*}(\alpha)$ serves as a good approximation to ${\bm{\beta}}^{*}={\bm{\beta}}^{*}_{0}$.
@@
Thanks to the convex program characterization of FPPE, the directional derivative $\mathop{}\!\mathrm{d}{\bm{\beta}}^{*}(\alpha)$ has a closed-form expression under certain regularity conditions (the conditions are given in Thm. 1; the full proof is given in the appendix). We need a few notations for this expression. Define ${\bm{\delta}}_{\alpha}=\int\nabla f(\theta,{\bm{\beta}}^{*}_{\alpha})(s-s^{\prime})\mathop{}\!\mathrm{d}\theta.$ Let ${\bm{H}}_{\alpha}=\nabla_{{\bm{\bm{\beta}}}{\bm{\bm{\beta}}}}^{2}\int F(\theta,{\bm{\beta}}^{*}_{\alpha})s_{\alpha}\mathop{}\!\mathrm{d}\theta$ be the Hessian matrix in the market ${\mathscr{M}}_{\alpha}$ and ${\bm{P}}_{\alpha}={\operatorname*{Diag}}(1({\bm{\beta}}^{*}_{\alpha,i}<1))$. Then, under the regularity conditions given in Thm. 1 below,
$\mathop{}\!\mathrm{d}{\bm{\beta}}^{*}(\alpha)=-({\bm{P}}_{\alpha}{\bm{H}}_{\alpha}{\bm{P}}_{\alpha})^{\dagger}{\bm{\delta}}_{\alpha}\;.$ (6)
We present a heuristic derivation in App. G. Given the closed-form expression of $\mathop{}\!\mathrm{d}{\bm{\beta}}^{*}(\alpha)$, we define the following debiased pacing multiplier
$\widetilde{{\bm{\beta}}}^{*}={\bm{\beta}}^{*}_{\alpha}+\alpha\cdot(-(\bm{P}_{\alpha}{\bm{H}}_{\alpha}{\bm{P}}_{\alpha})^{\dagger}{\bm{\delta}}_{\alpha})\;.$ (7)
@@
Suppose that in the market ${\mathscr{M}}_{0}$ conditions (SMO). and (SCS). hold, and assume that ${\bm{\bm{\beta}}}\mapsto\nabla^{2}\int F(\theta,{\bm{\bm{\beta}}})s^{\prime}\mathop{}\!\mathrm{d}\theta$ is twice continuously differentiable in a neighborhood of ${\bm{\beta}}^{*}$. Then the directional derivative $\mathop{}\!\mathrm{d}{\bm{\beta}}^{*}({\cdot})$ is well-defined in a neighborhood of zero, and is given by Eq. 6. Moreover, as $\alpha\downarrow 0$,
$\|\widetilde{{\bm{\beta}}}^{*}-{\bm{\beta}}^{*}\|_{2}=o(\alpha)\;.$
@@
The proof is given in Sec. B.1. Thm. 1 indicates that the debiased surrogate $\widetilde{{\bm{\beta}}}^{*}$ removes first-order bias caused by contamination. The limit pacing multipliers ${\bm{\beta}}^{*}_{\alpha}$ of the contaminated market ${\mathscr{M}}_{\alpha}$ will have bias of order ${\bm{\beta}}^{*}_{\alpha}-{\bm{\beta}}^{*}=\Theta(\alpha)$. In contrast, Thm. 1 shows that the debiased surrogate only incurs a bias of order $o(\alpha)$.
4.2 The Estimator
@@
In this section we introduce a plug-in estimator for the debiased surrogate $\widetilde{{\bm{\beta}}}^{*}$ and introduce a consistency theorem. The next section discusses constructing confidence intervals.
@@
To estimate $\mathop{}\!\mathrm{d}{\bm{\beta}}^{*}(\alpha)$ in Eq. 6 we need estimates of its three components: the Hessian ${\bm{H}}_{\alpha}=\nabla_{{\bm{\bm{\beta}}}{\bm{\bm{\beta}}}}^{2}\int F(\theta,{\bm{\beta}}^{*}_{\alpha})s_{\alpha}\mathop{}\!\mathrm{d}\theta$, the diagonal matrix ${\bm{P}}_{\alpha}$ and the vector ${\bm{\delta}}_{\alpha}=\int\nabla f(\theta,{\bm{\beta}}^{*}_{\alpha})(s-s^{\prime})\mathop{}\!\mathrm{d}\theta$.
@@
The Hessian. For simplicity in our theoretical results, we will simply assume a generic Hessian estimator $\widehat{\bm{H}}_{\alpha}$ such that for some $\eta_{t}\downarrow 0$ we have $\widehat{\bm{H}}_{\alpha}-{\bm{H}}_{\alpha}=O_{p}(\eta_{t})$. Hong et al. (2015) discuss the estimation of the derivative in detail. Different kinds of statistical guarantees require different rate conditions on $\eta_{t}$; see Theorems 2 and 3. We then introduce two Hessian estimators: one is applicable for general FPPE, while the other requires an extra market regularity condition. The first Hessian estimator is the finite difference method. Let $\bm{e}_{i},\bm{e}_{j}$ be basis vectors and $\varepsilon_{t}$ be a step-size. Then the estimator is
$\widehat{\bm{H}}_{\alpha}[i,j]=[H_{\alpha,t}({\bm{\beta}}^{\gamma}_{++})-H_{\alpha,t}({\bm{\beta}}^{\gamma}_{+-})-H_{\alpha,t}({\bm{\beta}}^{\gamma}_{-+})+H_{\alpha,t}({\bm{\beta}}^{\gamma}_{--})]/(4\varepsilon_{t}^{2})\;,$
@@
where ${\bm{\beta}}^{\gamma}_{\pm\pm}={\bm{\beta}}^{\gamma}_{\alpha}\pm\bm{e}_{i}\varepsilon_{t}\pm\bm{e}_{j}\varepsilon_{t}$, and $H_{\alpha,t}({\bm{\beta}})=\frac{1}{t}{\sum_{\tau=1}^{t}}F({\theta^{\tau}},{\bm{\beta}})$, with $\{{\theta^{\tau}}\}_{\tau}$ being the items in ^$\widehat{\mathscr{M}}_{\alpha}$. In practice, a diagonal approximation of the Hessian suffices. The second method relies on an additional regularity condition, in which case we derive a simplified formula for the Hessian, thereby enabling a simpler estimation procedure (see Thm. 3).
@@
The vector ${\bm{\delta}}_{\alpha}$. Let $g$ be the Radon-Nikodym ratio $g(\theta)=(\mathop{}\!\mathrm{d}(s-s^{\prime})/\mathop{}\!\mathrm{d}s_{\alpha})(\theta)=\frac{1}{1-\alpha}1_{\Theta_{\mathsf{good}}}(\theta)-\frac{1}{\alpha}1_{\Theta_{\mathsf{bad}}}(\theta)$. With the ratio $g$, the true vector ${\bm{\delta}}_{\alpha}$ can be written as ${\bm{\delta}}_{\alpha}=\int g(\theta)\nabla f(\theta,{\bm{\beta}}^{*}_{\alpha})s_{\alpha}(\theta)\mathop{}\!\mathrm{d}\theta$, which is easy to estimate given i.i.d. draws from $s_{\alpha}$. In particular, our estimator is then $\widehat{\bm{\delta}}_{\alpha}=\frac{1}{t}{\sum_{\tau=1}^{t}}g({\theta^{\tau}})\bm{\mu}^{\tau}\;.$ Here $\bm{\mu}^{\tau}=[x_{1}^{\tau}v_{1}^{\tau},\dots,x_{n}^{\tau}v_{n}^{\tau}]^{\scriptscriptstyle\mathsf{T}}$ is a subgradient of $f({\theta^{\tau}},{\bm{\beta}}^{\gamma}_{\alpha})$ w.r.t. ${\bm{\beta}}$.
@@
The diagonal matrix ${\bm{P}}_{\alpha}$. Recall ${\bm{P}}_{\alpha}={\operatorname*{Diag}}(1(\beta^{*}_{\alpha,i}<1))$. So a natural estimator is $\widehat{\bm{P}}_{\alpha}={\operatorname*{Diag}}(1(\beta^{\gamma}_{\alpha,i}<1-\iota_{t}))$, where the slackness $\iota_{t}\asymp\frac{1}{\sqrt{t}}$.
@@
With all three components estimated, using ${\bm{\beta}}^{\gamma}_{\alpha}$ is the pacing multiplier in the market ^$\widehat{\mathscr{M}}_{\alpha}$ we define the plug-in estimator for $\mathop{}\!\mathrm{d}{\bm{\beta}}^{*}(\alpha)$ in Eq. 6 as
$\widehat{\bm{\beta}}={\bm{\beta}}^{\gamma}_{\alpha}-\alpha\,{\cdot}\,(\widehat{\bm{P}}_{\alpha}\widehat{\bm{H}}_{\alpha}\widehat{\bm{P}}_{\alpha})^{\dagger}\widehat{\bm{\delta}}_{\alpha}\;.$ (8)
@@
4.3 Asymptotic Normality and Inference
We present two asymptotic normality results. In the first result, we require a stronger condition on the Hessian error rate $\eta_{t}$. In particular, as will be shown in Thm. 3, the rate condition $\eta_{t}=o(1/\sqrt{t})$ is sufficient for normality. One could use a separate large historical dataset to obtain a good estimate of the Hessian matrix. In the second result, we impose an additional condition on market structure which simplifies the Hessian expression and facilitates efficient Hessian estimation. To describe the additional market structure, we define the gap between the highest and the second-highest bid for an item $\theta$ under pacing ${\bm{\beta}}$ by ${\mathsf{{bidgap}}}({\bm{\beta}},\theta)=\max\{{\beta_{i}}{v_{i}(\theta)}\}-\operatorname{secondmax}\{{\beta_{i}}{v_{i}(\theta)}\}\;,$ where $\operatorname{secondmax}$ is the second-highest entry potentially equal to the highest; e.g., $\operatornamewithlimits{secondmax}([1,1,2])=1$. When there is a tie for an item $\theta$ under pacing ${\bm{\beta}}$, we have ${\mathsf{{bidgap}}}({\bm{\beta}},\theta)=0$. When there is no tie for an item $\theta$, the gap ${\mathsf{{bidgap}}}({\bm{\beta}},\theta)$ is strictly positive.
For any $g:\Theta\to{\mathbb{R}}$, let ${\mathbb{E}}_{\alpha}[g]=\int gs_{\alpha}\mathop{}\!\mathrm{d}\theta$ and ${{Cov}}_{\alpha}(g)={\mathbb{E}}_{\alpha}[(g-{\mathbb{E}}_{\alpha}[g])(g-{\mathbb{E}}_{\alpha}[g])^{\scriptscriptstyle\mathsf{T}}]$. Recall $\widetilde{{\bm{\beta}}}^{*}$ is the debiased surrogate in Eq. 7 and $\widehat{\bm{\beta}}$ is its estimator defined in Eq. 8.
We need to introduce a few more notations to describe the normality results. First, let $\bm{d}_{\alpha}=-({\bm{P}}_{\alpha}{\bm{H}}_{\alpha}{\bm{P}}_{\alpha})^{\dagger}\nabla f({\cdot},{\bm{\beta}}^{*}_{\alpha})\;$. As mentioned previously, the pacing multipliers in the contaminated market converge to the limit counterpart and have the representation
$\sqrt{t}({\bm{\beta}}^{\gamma}_{\alpha}-{\bm{\beta}}^{*}_{\alpha})=\frac{1}{\sqrt{t}}{\sum_{\tau=1}^{t}}(\bm{d}_{\alpha}({\theta^{\tau}})-{\mathbb{E}}_{\alpha}[\bm{d}_{\alpha}({\theta^{\tau}})])+o_{p}(1)\;.$
In the statistics literature, the function $\bm{d}_{\alpha}({\cdot})-{\mathbb{E}}[\bm{d}_{\alpha}]$ is called the influence function (van der Vaart, 2000). For our debiased estimators, we need the following (uncentered) influence functions.
$\bm{d}_{1}(\theta)=\big{(}\tfrac{1}{1-\alpha}1_{\Theta_{\mathsf{good}}}(\theta)\big{)}\bm{d}_{\alpha}(\theta)\;,\;\bm{d}_{2}(\theta)=\bm{d}_{1}(\theta)-2\alpha{\operatorname*{Diag}}(\beta^{*}_{\alpha,i}\delta_{\alpha,i}b_{i}^{{-1}})\bm{d}_{\alpha}(\theta)\;.$
Theorem 3.
Let (SCS). and (SMO). hold in ${\mathscr{M}}_{\alpha}$.